# ChatTS: Aligning Time Series with LLMs via Synthetic Data for Enhanced Understanding and Reasoning

### v2

@@ Abstract.
Understanding time series is crucial for its application in real-world scenarios. Recently, large language models (LLMs) have been increasingly applied to time series tasks, leveraging their strong language capabilities to enhance various applications. However, research on multimodal LLMs (MLLMs) for time series understanding and reasoning remains limited, primarily due to the scarcity of high-quality datasets that align time series with textual information. This paper introduces ChatTS, a novel MLLM designed for time series analysis. ChatTS treats time series as a modality, similar to how vision MLLMs process images, enabling it to perform both understanding and reasoning with time series. To address the scarcity of training data, we propose an attribute-based method for generating synthetic time series with detailed attribute descriptions. We further introduce Time Series Evol-Instruct, a novel approach that generates diverse time series Q&As, enhancing the model’s reasoning capabilities. To the best of our knowledge, ChatTS is the first TS-MLLM that takes multivariate time series as input for understanding and reasoning, which is fine-tuned exclusively on synthetic datasets. We evaluate its performance using benchmark datasets with real-world data, including six alignment tasks and four reasoning tasks. Our results show that ChatTS significantly outperforms existing vision-based MLLMs (e.g., GPT-4o) and text/agent-based LLMs, achieving a 46.0% improvement in alignment tasks and a 25.8% improvement in reasoning tasks. We have open-sourced the source code, model checkpoint and datasets at https://github.com/NetManAIOps/ChatTS.
@@ 1. Introduction
Multimodal large language models (MLLMs) have recently achieved significant progress in vision-language tasks, showing exceptional performance even in scenarios requiring complex understanding and reasoning (Liu et al., 2024a; Bai et al., 2023; Li et al., 2023b; Yin et al., 2024). However, this success has not been replicated in the time series domain. Even though some studies have attempted to integrate LLMs with time series, such as TimeLLM (Jin et al., 2023), they usually only focus on specific classical time series tasks (e.g., forecasting) rather than complex understanding and reasoning. Moreover, recent studies indicate that LLMs still struggle with zero-shot reasoning about time series (Merrill et al., 2024). This is particularly significant because time series analysis, widely applied in domains such as electricity (Wolde-Rufael, 2006), healthcare (Penfold and Zhang, 2013), traffic (Li et al., 2015), weather (Lim and Zohren, 2021), and finance (Sezer et al., 2020), frequently requires understanding and reasoning about data to identify patterns, trends, and events over time. Therefore, the ability to reason using both text and time series data is a critical capability for MLLMs, enabling them to support human decision-making by providing natural language explanations that align with human logic. Figure 1 illustrates such an example in an AIOps (Zhong et al., 2023) scenario where understanding and reasoning about multivariate system monitoring time series are achieved through natural language dialogue, thereby improving the diagnostic and troubleshooting process.
Figure 1. Example of an AIOps application of time series-related dialogue. The LLM answers the user’s questions based on the input time series.
Existing LLM-based methods for time series understanding and reasoning can be broadly categorized into text-based, vision-based, and agent-based approaches. Text-based methods directly use LLMs by structuring historical observations as raw text (Alnegheimish et al., 2024). However, these methods are often constrained by the limitation of prompt length and generally perform poorly in understanding the global features of time series compared to vision-based methods. Vision-based methods utilize vision MLLMs, which accept plot figures of time series data (Merrill et al., 2024), such as GPT-4o (gpt, 2024) or Qwen-VL (Bai et al., 2023). While these methods can better capture global features, they are limited by the resolution of the plotted figures and face challenges in accurately interpreting detailed fluctuations. Recent works also demonstrate how agents can leverage time series analysis tools to interact with LLMs (Zhou et al., 2023a; Wang et al., 2024). However, the ability of agents to understand time series is restricted by the functionality of the tools.
@@
However, developing TS-MLLMs with effective time series understanding and reasoning ability faces several core challenges. First, multimodal time series data, especially language-time series pair data, is extremely scarce (Merrill et al., 2024; Chow et al., 2024; Jin et al., 2024). Unlike modalities such as images and audio, almost no research focuses on the language alignment of time series. As a result, there is a significant lack of time-series + text data, which makes the construction of time-series dialogue and reasoning datasets challenging. This is fundamental for TS-MLLMs to develop temporal understanding and reasoning capabilities. Second, time-series data contains abundant shape and numerical attributes (i.e., the types of local fluctuations and their amplitudes). Therefore, a diverse range of text is needed to comprehensively describe these attributes while ensuring accuracy to achieve effective alignment. Third, real-world time-series data are usually variable in length, multivariate, and of uncertain quantity. The correlations among MTS are often a focus of attention (as illustrated in Figure 1). In MLLMs for other modalities, such as images, few methods emphasize the relationships between multiple samples. However, such relationships are indispensable for understanding and reasoning about time series. Finally, there is a lack of evaluation data and methods for TS-MLLMs. Developing comprehensive and reasonable datasets and methodologies to evaluate their performance is necessary.
@@
  • We propose a context-aware TS-MLLM, ChatTS, designed for variable-length, multivariate time series input and trained using the generated synthetic data. To the best of our knowledge, ChatTS is the first TS-MLLM with multivariate time series as input for understanding and reasoning.
@@
  • The model, source code, and evaluation datasets have been released at: https://github.com/NetManAIOps/ChatTS for reproduction and further research.
@@ 2.2. Existing Methods
Although mainstream LLMs currently do not support the direct input of time series modality data, time series information can be provided to LLMs through alternative methods to do simple time series understanding and reasoning tasks as shown in Figure 2. Existing approaches can be broadly categorized into text-based, vision-based, and agent-based, each with distinct limitations.
@@ 3.1. Overview
  • Model Training (Section 3.5): A large-scale pretraining and a supervised fine-tuning are conducted to perform language alignment and improve time series-related reasoning ability.
@@ 3.2. Attribute-Based Time Series Generator
Diverse time series and precise, detailed textual attribute descriptions are essential to achieve accurate time series language alignment. Time series have rich pattern attributes, which can be roughly categorized into trend, periodicity, and remainder (RB, 1990). Much existing research on the generation of time series (Fu et al., 2024; Fons et al., 2024) also adopts similar approaches to classify these attributes. Therefore, following existing studies, we classify time series attributes into four major categories, Trend, Periodicity, Noise, and Local Fluctuation, to construct the corresponding attribute set for time series.
Based on this, we propose an attribute selector and an attribute-based time series generator that produces synthetic time series data (see Figure 4). First, we define an “All Attribute Set”, which includes many specific attributes under different attribute categories. The All Attribute Set includes 4 types of Trend, 7 types of Seasonality, 3 types of Noise, and 19 types of local fluctuations. The complete list can be found in the source code. Different attributes within the same category can be combined. For example, a time series can contain multiple segments of trends and several local fluctuations. This allows the generator to generate an unlimited variety of attribute combinations theoretically. We also introduced a GPT Selector. Specifically, when generating an attribute set for time series, we randomly sample a metric from a large “Metric Set” that contains 567 predefined metric names [^1] The list of the metric names can be found in the source code. from real-world scenarios and use GPT to choose a attribute subset from the all attribute set, based on the actual physical meaning of the metric and the predefined scenario. This helps align time series with real-world physical meanings.
@@ 3.4.2. Value-Preserved Time Series Normalization
The numerical features of time series are essential, as real-world applications often involve specific numerical queries (e.g., asking for the maximum CPU utilization). However, normalization of time series data can lead to losing original numerical information. To address this, we introduce a value-preserved time series normalization scheme (as shown in Figure 7). First, we apply standard min-max normalization (0-1 scaling) to each time series array. Then, for each time series, we include the normalization parameters-“Value Scaling” (the scaling factor during normalization) and “Value Offset” (the offset applied during normalization)—in the text as part of a prompt. This approach leverages the numerical understanding capabilities of LLMs, enabling us to normalize time series features while preserving the original numerical information.
@@ 3.5. Model Training
ChatTS is trained based on QWen2.5-14B-Instruct (Yang et al., 2024) [^2] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct, with a two-stage fine-tuning process: large-scale alignment pretraining and supervised fine-tuning (SFT). Table 1 shows the datasets we use during training.
3.5.1. Large-Scale Alignment Pretraining
In the first stage, we perform large-scale alignment pretraining using the attribute-based synthetic time series data. During the alignment stage, we created three datasets for large-scale pretraining based on a series of manually designed templates (which can be found in the supplemental material). The UTS dataset includes tasks for basic attribute descriptions of univariate time series. The MTS-Shape dataset consists of multivariate data with overall trend correlations designed to enhance the model’s ability to analyze multivariate correlations. The MTS-Local dataset contains multivariate data with correlated local fluctuations, aiming to improve the model’s capability in analyzing local features of multivariate data. This stage aims to establish an initial alignment between the text and time series modalities within the LLM, enabling it to align textual descriptions with attributes in time series data effectively. With a large amount of synthetic data, the model acquires foundational multimodal alignment capabilities, which are essential for alignment and reasoning tasks.
@@ 3.5.2. Supervised Fine-Tuning
In the second stage, we use SFT to develop the LLM’s ability to perform complex question-answering and reasoning tasks. This stage utilizes two main types of training data: the Q&A datasets generated with TSEvol, designed to enhance the model’s QA and reasoning ability about time series, and an instruction-following (IF) dataset, constructed based on a series of predefined templates, designed to enhance the model’s ability to follow specific response formats. Together, these datasets train the multimodal LLM to respond accurately to time series-specific queries and follow task instructions, strengthening its capacity for complex, context-driven question-answering and reasoning tasks.
@@ 4. Evaluation
  • RQ3. Is the time series modalilty in ChatTS truly useful?
  • RQ4. Are attribute-based data generation and Time Series Evol-Instruct effective?
  • RQ5. Does ChatTS, with its native time-series multimodal capabilities, have advantages over agent-based methods?
@@ 4.1.1. Evaluation Tasks
To comprehensively evaluate the model’s performance, we set two categories of evaluation tasks: alignment tasks and reasoning tasks, following the general evaluation methods of other multimodal LLMs (Liu et al., 2024a; Luo et al., 2024). For each type of evaluation task, we designed a series of subtasks based on existing work. Some example QAs are shown in Figure 8 (more details can be found in the code and evaluation datasets). Specific tasks that rely heavily on domain-specific knowledge (e.g., classification and etiological reasoning) were excluded due to the lack of high-quality datasets that provide sufficient background information. Therefore, we primarily focused on the following tasks:
@@
More details about the evaluation tasks can be found in the evaluation dataset.
@@ B Alignment (Trend, Season, Noise, Local, Correlation, Cluster), Reasoning (Inductive) 1,616
Dataset A includes real-world time series data collected from multiple domains, including AIOps (Li et al., 2022b), weather [^3] https://www.bgc-jena.mpg.de/wetter/, the NAB (Numenta Anomaly Benchmark) (Ahmad et al., 2017), and Oracle system metrics (Li et al., 2022a). We manually label and collect a total of 525 questions, including both alignment tasks and reasoning tasks.
@@
MCQ2 (Merrill et al., 2024) is an open-source dataset [^4] https://github.com/behavioral-data/TSandLanguage that includes comparison reasoning tasks. The questions, answers, and time series in this dataset are all generated by LLMs. We did not use the etiological reasoning and forecasting datasets as they are not aligned with our evaluation settings. Furthermore, (Merrill et al., 2024) suggests that the settings of the MCQ1 dataset are unsuitable for evaluating the performance of time series reasoning, so we also did not adopt it. Considering the inference cost, we randomly sampled 100 questions.
@@ 4.1.4. Baselines
  • Agent-Based: These methods employ the ReAct (Yao et al., 2022) framework to interact with multiple tools to analyze the time series. The tools used include single-point/range query, STL decomposition, anomaly detection (autoregression AD in adtk [^5] https://github.com/arundo/adtk), classification (Rocket (Dempster et al., 2020)), and Pearson correlation. We choose GPT-4o/GPT-4o-mini as LLMs for the agent baselines. We also conducted additional experiments to explore further the capabilities of agent-based methods (Section 4.6), which studies the impact of tool accuracy.

### v3

@@ Abstract.
Understanding time series is crucial for its application in real-world scenarios. Recently, large language models (LLMs) have been increasingly applied to time series tasks, leveraging their strong language capabilities to enhance various applications. However, research on multimodal LLMs (MLLMs) for time series understanding and reasoning remains limited, primarily due to the scarcity of high-quality datasets that align time series with textual information. This paper introduces ChatTS, a novel MLLM designed for time series analysis. ChatTS treats time series as a modality, similar to how vision MLLMs process images, enabling it to perform both understanding and reasoning with time series. To address the scarcity of training data, we propose an attribute-based method for generating synthetic time series with detailed attribute descriptions. We further introduce Time Series Evol-Instruct, a novel approach that generates diverse time series Q&As, enhancing the model’s reasoning capabilities. To the best of our knowledge, ChatTS is the first MLLM that takes multivariate time series as input for understanding and reasoning, which is fine-tuned exclusively on synthetic datasets. We evaluate its performance using benchmark datasets with real-world data, including six alignment tasks and four reasoning tasks. Our results show that ChatTS significantly outperforms existing vision-based MLLMs (e.g., GPT-4o) and text/agent-based LLMs, achieving a 46.0% improvement in alignment tasks and a 25.8% improvement in reasoning tasks. We have open-sourced the source code, model checkpoint and datasets at https://github.com/NetManAIOps/ChatTS.
@@ 1. Introduction
Multimodal large language models (MLLMs) have recently achieved significant progress in vision-language tasks, showing exceptional performance even in scenarios requiring complex understanding and reasoning (Liu et al., 2024a; Bai et al., 2023; Li et al., 2023b; Yin et al., 2024). However, this success has not been replicated in the time series domain. Even though some studies have attempted to integrate LLMs with time series, such as TimeLLM (Jin et al., 2023), they usually only focus on specific classical time series tasks (e.g., forecasting) rather than understanding, reasoning, and dialogue based on time series attributes, as well as integrating into existing LLM workflows. Moreover, recent studies indicate that LLMs still struggle with zero-shot reasoning about time series (Merrill et al., 2024). This is particularly significant because time series analysis, widely applied in domains such as electricity (Wolde-Rufael, 2006), healthcare (Penfold and Zhang, 2013), traffic (Li et al., 2015), weather (Lim and Zohren, 2021), and finance (Sezer et al., 2020), frequently requires understanding and reasoning about time series patterns. Therefore, the ability to reason using both text and time series data is a critical capability for MLLMs, enabling them to support human decision-making by providing natural language explanations that align with human logic. Figure 1 illustrates such an example in an AIOps (Zhong et al., 2023) scenario where understanding and reasoning about multivariate system monitoring time series are achieved through natural language dialogue, thereby improving the diagnostic and troubleshooting process.
Figure 1. Example of an AIOps application of time series-related dialogue.
Existing LLM-based methods for understanding and reasoning about time series attributes can be broadly categorized into text-based, vision-based, and agent-based approaches. Text-based methods directly use LLMs by structuring historical observations as raw text (Alnegheimish et al., 2024). However, these methods are often constrained by the limitation of prompt length and generally perform poorly in understanding the global features of time series compared to vision-based methods. Vision-based methods utilize vision MLLMs, which accept plot figures of time series data (Merrill et al., 2024), such as GPT-4o (gpt, 2024) or Qwen-VL (Bai et al., 2023). While these methods can better capture global features, they are limited by the resolution of the plotted figures and face challenges in accurately interpreting the details. Recent works also show how agents can leverage time series analysis tools to interact with LLMs (Zhou et al., 2023a; Wang et al., 2024). However, the ability of agents to understand time series is restricted by the functionality of the tools.
@@
However, developing TS-MLLMs with effective understanding and reasoning ability for time series attributes faces several core challenges. First, multimodal time series data, especially language-time series pair data, is extremely scarce (Merrill et al., 2024; Chow et al., 2024; Jin et al., 2024). Unlike modalities such as images and audio, almost no research focuses on the language alignment of time series. As a result, there is a significant lack of time-series + text data, which makes the construction of time-series dialogue and reasoning datasets challenging. This is fundamental for TS-MLLMs to develop temporal understanding and reasoning capabilities. Second, time-series data contains abundant shape and numerical attributes (i.e., the types of local fluctuations and their amplitudes). Therefore, a diverse range of text is needed to comprehensively describe these attributes while ensuring accuracy to achieve effective alignment. Third, real-world time-series data are usually variable in length, multivariate, and of uncertain quantity. The correlations among MTS are often a focus of attention (as illustrated in Figure 1). In MLLMs for other modalities, such as images, few methods emphasize the relationships between multiple samples. However, such relationships are indispensable for understanding and reasoning about time series. Finally, there is a lack of evaluation data and methods for TS-MLLMs. Developing comprehensive and reasonable datasets and methodologies to evaluate their performance is necessary.
@@
  • We propose a context-aware TS-MLLM, ChatTS, designed for variable-length, multivariate time series input and trained using the generated synthetic data. To the best of our knowledge, ChatTS is the first TS-MLLM with multivariate time series as input for understanding and reasoning about time series attributes.
@@
  • We have open-sourced the model, source code, and evaluation datasets to support future research: https://github.com/NetManAIOps/ChatTS.
@@ 2.2. Existing Methods
Although mainstream LLMs currently do not support the direct input of time series modality data, time series information can be provided to LLMs through alternative methods to do simple understanding and reasoning about time series attributes, as shown in Figure 2. Existing approaches can be broadly categorized into text-based, vision-based, and agent-based, each with distinct limitations.
@@ 3.1. Overview
  • Model Training (Section 3.5): A large-scale training and a SFT are conducted to perform language alignment and improve time series-related reasoning ability.
As shown in Figure 3, the framework in ChatTS integrates synthetic data generation and model training into a pipeline that ensures effective time series attributes understanding and reasoning with only synthetic data. First, building on the attribute-based time series generation and TSEvol described in Sections 3.2 and 3.3, the pipeline generates synthetic data that captures intricate numerical and textual information for effective multimodal alignment. This data is then used to train the model (Section 3.4), where the context-aware time series encoder preserves the time series values while aligning attributes with textual semantics accurately. Finally, by alignment training and SFT (Section 3.5), ChatTS achieves precise alignment between time series encoding and text embeddings, along with enhanced reasoning capabilities.
@@ 3.2. Attribute-Based Time Series Generator
Diverse time series and precise, detailed textual attribute descriptions are essential to achieve accurate time series language alignment. Time series have rich pattern attributes, which can be roughly categorized into trend, periodicity, and remainder (RB, 1990; He et al., 2023). Much existing research on the generation of time series (Fu et al., 2024; Fons et al., 2024) also adopts similar approaches to classify these attributes. Therefore, following existing studies, we classify time series attributes into four major categories, Trend, Periodicity, Noise, and Local Fluctuation, to construct the corresponding attribute set for time series.
Based on this, we propose an attribute selector and an attribute-based time series generator that produces synthetic time series data (see Figure 4). First, we define an “All Attribute Set”, which includes many specific attributes under different attribute categories. The All Attribute Set includes 4 types of Trend, 7 types of Seasonality, 3 types of Noise, and 19 types of local fluctuations. The complete list can be found in the source code. Different attributes within the same category can be combined. A time series can include multiple segments of trends and several local fluctuations by combining the same type of attributes (see Figure 4). Additionally, by combining sine waves, we can generate a diverse range of periodic fluctuation patterns. Therefore, the proposed time series generator can theoretically generate an infinite number of different time series, ensuring the richness of attributes. We also introduced a GPT Selector. Specifically, when generating an attribute set for time series, we randomly sample a metric from a large “Metric Set” that contains 567 predefined metric names from real-world scenarios and use GPT to choose a attribute subset from the all attribute set, based on the actual physical meaning of the metric and the predefined scenario. This helps align time series with real-world physical meanings.
@@ 3.4.2. Value-Preserved Time Series Normalization
The numerical features of time series are essential, as real-world applications often involve specific numerical queries (e.g., asking for the maximum CPU utilization). However, normalization of time series data can lead to losing original numerical information. To address this, we introduce a value-preserved time series normalization scheme (as shown in Figure 7). First, we apply standard min-max normalization (0-1 scaling) to each time series array. Then, for each time series, we include the normalization parameters-“Value Scaling” (the scaling factor during normalization) and “Value Offset” (the offset applied during normalization)—in the text as part of a prompt. This approach leverages the numerical understanding capabilities of LLMs, enabling us to normalize time series features while preserving the original numerical information. To further enhance numerical understanding, numerical tasks are included in the training dataset (see Section 3.5).
@@ 3.5. Model Training
ChatTS is trained based on QWen2.5-14B-Instruct (Yang et al., 2024) [^1] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct, with a two-stage fine-tuning process: large-scale alignment training and supervised fine-tuning (SFT). Table 1 shows the datasets we use during training.
3.5.1. Large-Scale Alignment Training
In the first stage, we perform large-scale alignment training using the attribute-based synthetic time series data to establish an initial alignment between the text and time series modalities within the LLM. This stage enables ChatTS to align textual descriptions with time series attributes effectively. During the alignment stage, we created three datasets for large-scale training based on a series of manually designed templates and LLM refinement. The UTS dataset includes tasks for basic attribute descriptions of univariate time series (both global and local attribute tasks are included). The MTS-Shape dataset consists of multivariate data with global trend correlations designed to enhance the model’s ability to analyze multivariate correlations. The MTS-Local dataset contains multivariate data with correlated local fluctuations, aiming to improve the model’s capability in analyzing local features of multivariate data. Given MTS’s more complex feature combinations, we set the training data size for MTS and UTS at an approximately 2:1 ratio. We conduct a dataset scaling study in Section 4.5 to investigate the impact of training dataset size.
@@ 3.5.2. Supervised Fine-Tuning
In the second stage, we use SFT to develop the LLM’s ability to perform complex question-answering and reasoning tasks. This stage utilizes two main types of training data: the datasets generated with TSEvol, designed to enhance the model’s question answering and reasoning ability about time series, and an instruction-following (IF) dataset, constructed based on a series of predefined templates, designed to enhance the model’s ability to follow specific response formats. For TSEvol, we used the dataset from alignment training along with LLM-generated QAs as the seed data. Together, these datasets train the multimodal LLM to respond accurately to time series-specific queries and follow task instructions, strengthening its capacity for complex, context-driven question-answering and reasoning tasks. In both Alignment and SFT stages, we enhance ChatTS’s numerical capabilities through a series of numerical tasks. Specifically, we explicitly train the model to learn various aspects, such as maximum/minimum values, segmented averages, local features (e.g., spike positions and amplitudes), seasonality and trend amplitudes, and raw numerical values at individual time points. The numerical evaluation metrics in our experimental results further demonstrate ChatTS’s strong performance in time series numerical analysis.
3.5.3. Training Settings
We use QA pairs as the data format for both training stages. During alignment training, we mixed in a small amount of IF data and found that this mitigates the decline in the model’s IF ability. In the SFT stage, we mixed 30% of the alignment training dataset to reduce overfitting. The training dataset includes time series with lengths ranging from 64 to 1024 to ensure that ChatTS can handle varying time series lengths. Full-parameter SFT is used for ChatTS with DeepSpeed (dee, 2020) and LLaMA-Factory (Zheng et al., 2024), with Qwen2.5-14B-Instruct (Yang et al., 2024; qwe, 2024) as the base model. Inference for both Qwen and ChatTS is also conducted with DeepSpeed.
@@ 4. Evaluation
  • RQ3. Are attribute-based data and TSEvol effective?
  • RQ4. How does the training set size affect model performance?
  • RQ5. Is the time series modalilty in ChatTS truly useful?
  • RQ6. Does ChatTS, with its native time-series multimodal capabilities, have advantages over agent-based methods?
@@ 4.1.1. Evaluation Tasks
To comprehensively evaluate the model’s performance, we set two categories of evaluation tasks: alignment tasks and reasoning tasks, following the general evaluation methods of multimodal LLMs (Liu et al., 2024a; Luo et al., 2024; Cai et al., 2024). For each type of evaluation task, we designed a series of subtasks based on existing work. Some example QAs are shown in Figure 8 (more details can be found in the source code). Specific tasks that rely heavily on domain-specific knowledge (e.g., classification and etiological reasoning) were excluded due to the lack of high-quality datasets that provide sufficient background information. Therefore, we primarily focused on the following tasks:
@@
More details about the evaluation tasks can be found in the source code and the evaluation dataset.
@@ B Alignment (Trend, Season, Noise, Local, Correlation, Cluster), Reasoning (Inductive) 1,616
Dataset A includes real-world time series data collected from multiple domains, including AIOps (Li et al., 2022b), weather (wea, 2023), the NAB (Numenta Anomaly Benchmark) (Ahmad et al., 2017), and Oracle system metrics (Li et al., 2022a). We manually label and collect a total of 525 questions, including both alignment tasks and reasoning tasks.
@@
MCQ2 (Merrill et al., 2024) is an open-source dataset (mcq, 2024) that includes comparison reasoning tasks. The questions, answers, and time series in this dataset are all generated by LLMs. We did not use the etiological reasoning and forecasting datasets as they are not aligned with our evaluation settings. Furthermore, (Merrill et al., 2024) suggests that the settings of the MCQ1 dataset are unsuitable for evaluating the performance of time series reasoning, so we also did not adopt it. Considering the inference cost, we randomly sampled 100 questions.
@@ 4.1.4. Baselines
  • Agent-Based: These methods employ the ReAct (Yao et al., 2022) framework to interact with multiple tools to analyze the time series. The tools used include single-point/range query, STL decomposition, anomaly detection (autoregression AD in adtk (adt, 2019)), and classification (Rocket (Dempster et al., 2020)) for UTS; trend/fluctuation correlation (based on Pearson correlation & rules), multivariate version of AD and classification for MTS. We choose GPT-4o/GPT-4o-mini for the agent. More details about the tools’ implementation can be found in the source code. We also conducted additional experiments to explore further the capabilities of agent-based methods (Section 4.7), which studies the impact of tool accuracy.